# Çukurova Üniversitesi Rektörlüğü 01250 Sarıçam, Adana, Turkey

### v1

@@ 2 The HGCAL Prototype, event reconstruction and simulation
The HGCAL prototype detector used in the beam test in 2018 consisted of a silicon sampling electromagnetic section (CE-E prototype) and a silicon sampling hadronic section (CE-H prototype). These were followed by the CALICE AHCAL prototype, which is a sampling calorimeter using scintillator tiles directly read out by SiPMs bib.calice-ahcal-2018tb-paper . The experimental setup is described in bib.cern-h2paper and is shown in Fig. 1. The basic detector unit of the CE-E and CE-H prototypes was a silicon module consisting of a 6-inch hexagonal silicon sensor and a printed circuit board (PCB) with the readout electronics, along with a baseplate for mechanical support. Each silicon sensor was subdivided into 128 hexagonal cells with an area of approximately 1.1 $cm^{2}$, and each cell was readout independently with signal amplitude and timing information being collected.
@@
For events collected, the digitized signal from each readout channel was converted into the corresponding number of minimum ionizing particles (MIPs). The MIP conversion factor for each channel is obtained by fitting the digitized signal obtained with 200 GeV muons with a Landau distribution convoluted with a Gaussian distribution. The maxima of the fitted function is the MIP calibration value for that cell.
The detector setup, along with the beamline elements, was simulated using a Geant4 simulation bib.GEANT4 . The beamline elements, starting from the production target T2 up to the front of the HGCAL prototype, were simulated using the G4Beamline simulation framework implemented in Geant4 version 10.3. The calorimeter sections were simulated using Geant4 version 10.4.3. A schematic of the simulated HGCAL and AHCAL prototype detectors used in the beam test experiment is shown in Fig. 2. The details are summarized in Ref. bib.hgcal-2018-pions . The simulated signal is also converted into units of the number of MIPs using the same procedure as used for the data. The position of cells is defined in a right-handed Cartesian coordinate system with the z-axis parallel to the beam direction.
Figure 2: Simulated geometry of the CE-E, CE-H and AHCAL detector prototypes depicting the position of active layers and absorbing material.
A sample of five million charged pions with a flat energy distribution between 10 and 350 GeV were produced using the Geant4 toolkit implemented in the CMS software framework with FTFP_BERT_EMN physics list bib.g4validation . This sample is used to train the neural network models investigated for the energy reconstruction of hadron showers produced by charged pions.
@@ 3 Machine learning strategy and the dynamic reduction network
The highly granular nature of the HGCAL makes it very suitable for the application of machine learning (ML) techniques to reconstruct particle showers in the calorimeter. Compared to other calorimeter designs, the degree of information is significantly higher and ML techniques, which take as input all of the available information or a large subset of it, can be used to exploit the detailed information to reconstruct hadronic and electromagnetic showers.
@@
These showers have large event-by-event variations in the spatial distributions of energy deposits in the detector cells and are usually interspersed with channels without a measurable signal. Traditional ML architectures like multilayer perceptron neural networks (MLPs) bib.tensorflow2015-whitepaper and boosted decision trees (BDTs), expect fixed-size human-designed inputs features for reconstructing these showers, and it is difficult to design such features to efficiently capture these variations. Modern ML algorithms possess powerful feature-learning capabilities, and have demonstrated their ability to outperform other ML algorithms, which use human-designed input features, particularly when provided with low level detector information, like calorimeter hits ImageNet ; CNNNeutrino ; DNNMicroBoone ; CMSHtoAA . While image based approaches like convolutional neural networks (CNNs) do use low level detector information, the requirement of fixed size inputs necessitate zero-padding or truncation of the input collection. Additionally it is not possible to naturally represent information from multiple layers of detectors with varying granularity in square or rectangular images. Finally, these architectures tend to impose ordering on their inputs that is inappropriate for this permutation-invariant collection of calorimeter hits. A natural way to represent these sparse, high-dimensional, variable-geometry sets of input features is with graphs. In particular, we represent each pion object as a point cloud of hits in a high-dimensional latent space and then create graphs by drawing edges between nearby hits. Information is allowed to flow along the graph edges in GNNs as in CNNs, but with the advantage of, among others, flexibility in morphology.
@@
In electromagnetic showers, bremsstrahlung and pair-production processes dominate the energy loss until the shower components’ energy falls below the critical energy, typically of the order of 10 MeV. The fluctuations in the amount of energy that is undetected are small and do not affect significantly the energy resolution. However, in hadronic showers a large component of the energy is lost through excitation or disintegration of the absorber nuclei, and, since it is generally undetected, it is known as the ‘invisible energy’ and can amount to up to 40% of the total energy of a shower. Additionally charged and neutral pions produced in nuclear interactions leave different energy depositions, unless the calorimeter is specifically designed to be ‘compensating’. Fluctuations in both the production rate of neutral or charged pions, and in the energy lost as ‘invisible energy’ lead to significant event-by-event variations in the energy response of a calorimeter to incoming hadrons.
@@
In the conventional WS method used to determine the energy of an incoming hadron in a calorimeter that is divided into a few sampling layers (for example the 4 or 5 layers of the current CMS hadron calorimeter bib.cms-phase-one-upgrade ), where the first layer is usually an electromagnetic calorimeter. To reconstruct a hadronic shower, the energy deposition in each layer is added with a different weight that is determined either empirically or by simulation. This method does not fully account for the large event-by-event fluctuations in hadronic showers and requires fine tuning for different regions of the detector, as well as for any dependence on the energy of incident hadrons. This method works well for electromagnetic showers, which tend to be uniform, whereas for hadronic showers fluctuations limit its efficacy bib.hgcal-em-optimization . In the context of high granularity calorimeters, the method has been extended by using local energy densities as a measure of the electromagnetic component of hadronic showers as described in bib.calice-software-compensation .
@@
The HGCAL, with its fine longitudinal and transverse granularity, produces a highly detailed image of the longitudinal and lateral development of the shower. By using ML methods the information in this three-dimensional image can be used in a structured way to derive the energy and other properties of the incoming hadron. In addition, in the HGCAL, the arrival time of signal in each readout out channel will be available, adding another dimension to the shower reconstruction. The possibilities that this provides for event reconstruction have not been explored in this work.
@@
The DRN bib.drn-pointclouds builds upon dynamic graph convolution networks bib.drn with two modifications: firstly the graph operations are performed in a higher dimensional latent space, and secondly there is an additional graph clustering and pooling step used to iteratively accumulate information across the graph. A schematic of such a network used for the studies presented in this paper is shown in Fig. 3. A summary of the information flow from the inputs – the spatial coordinates of the hits and their energy measured in each detector readout channel – to the required target is as follows.
@@
    The input features, attributes of detector channels, are normalized using fixed values such that most of the values lie in the range from zero to one, with outliers allowed. The normalized input data are mapped onto a 64-dimensional latent space using a fully connected neural network, which is an MLP encoder with a depth of three layers that are connected via an exponential linear unit activation function.
@@
    High-level features are learnt using GNN techniques by repeating the following three steps twice:
@@
      Graph convolution step: the resulting graphs are processed using the “EdgeConv” operation bib.edgeconv , in which two layer deep fully connected networks are used to construct ’messages’ that are passed along the graph edges. These ’messages’ are aggregated at each node in the graph to form the new node coordinates.
@@
      Graph clustering and pooling: the graph convolution step is followed by graph clustering using the graclus clustering algorithm bib.graclus and a pooling step, in which the updated features are weighted by distance and the clusters are combined using a max-pooling strategy to form a single point. The graclus clustering only pairs together nodes and does not form arbitrarily large-sized clusters. Each pass through these three steps – the graph generation, graph convolutions, and graph clustering and pooling – halves the number of nodes in the graph.
@@
As described in Ref. bib.drn-pointclouds , the clustering and pooling steps are optimized in a way to maximize reduction in information passed on for the further processing, hence the network is called a ‘dynamic reduction network (DRN)’.
Figure 3: Schematic of the operation of the GNN based dynamic reduction network used in this analysis.
@@ 4 Energy Reconstruction of pion showers using DRN method
A sample of simulated hadron showers produced by negatively charged pions, uniformly sampled in the energy range from 10 to 350 GeV, is used to train the DRN for the reconstruction of the energy of incident pions. The hadron showers were simulated using the same Geant4 geometry and physics lists that were used for the simulation of events in the beam tests described above. We apply the same selection criteria to remove events where the particles loose a large fraction of their energy in the beamline and apply the same noise rejection thresholds that were used in the data bib.hgcal-2018-pions . The input to the DRN is a point cloud of the detector channels associated with a given hadronic shower. The attributes of each channel, used for training the model, are its position coordinates, $x$, $y$ and $z$, given in centimeters and the amplitude of the energy deposition in the sensor as either the number of MIPs or in terms of the effective energy in GeV, defined below.
As this is a sampling calorimeter, we need to infer from the signal in each sensor the energy deposited in the absorbers. In the prototype detector, the absorbers in CE-E, CE-H and AHCAL were different and hence energy scale factors need to be estimated separately for them. We estimate the energy deposited in the absorber that corresponds to the observed signal amplitude using a MIP-to-GeV conversion factor. For CE-E this factor is estimated with 50 GeV positrons. In CE-H and AHCAL the MIP-to-GeV conversion factor is estimated with 50 GeV pions that do not initiate a hadronic shower in CE-E. We obtain factors of approximately 10.5 MeV per MIP and 80 MeV per MIP for the electromagnetic and hadronic sections, respectively.
The signals in the detector cells, weighted by their corresponding MIP-to-GeV factor, are used to estimate of the amount of energy deposited in the absorber plates. These estimates of energy deposited in the absorber together with their positions are referred to as ’reconstructed hits’ or ’rechits’. Distributions of the number of rechits with a signal measured above the noise thresholds in CE-E, CE-H and AHCAL sections, and their $x$ and $y$ coordinates are shown in Fig. 4 (top left) and Fig. 4 (top right), respectively, for simulated 200 GeV $\pi^{-}$ data. The distributions of the energy estimated in units of GeV in the detector cells of the three detector sections are shown in Fig. 4 (bottom left). Fig. 4 (bottom right) shows the distribution of rechit energies along the depth of the detector (z coordinate).
@@
The energies of all rechits can be summed to obtain an estimate of the total energy of the particles; this sum is referred to as $E_{fix}$ in the following. However, since the response of a sampling calorimeter, such as the HGCAL prototype, to hadrons is not linear with the incident particle’s energy, $E_{fix}$ is a suboptimal estimate of the particle’s true energy. The WS method uses a $\chi^{2}-minimization$ method for a given incident particle energy, as discussed in bib.hgcal-2018-pions , to obtain weights to combine energies measured in CE-E, CE-H, and AHCAL sections. Although the linearity of the response can be restored and the resolution improved, this method does not account for event-by-event fluctuations in the hadronic showers.
@@
We have trained a DRN model with simulated hadronic showers generated with a flat incident energy distribution. The target in the ML training was the ratio of true energy of the particle to that measured using the fixed MIP-to-GeV weights ($E_{fix}$). In the training the learning rate was set to be 0.0001. We trained three different models with three different input features, the energy ($E$), the energy and the longitudinal position ($E,z$), and the energy and both the lateral and the longitudinal positions ($E,x,y,z$) of each detector channel. These models are trained until the value of loss function evaluated on the validation dataset stopped improving in subsequent epochs. The epoch numbers 110, 114 and 135 correspond to the minimum loss values for DRN(E), DRN(E,z), and DRN(E,x,y,z), respectively, and corresponding training weights are used for evaluating the performance of pion energy reconstruction in the following. Approximately 20% of the generated events were reserved in a separate dataset for validation of the models using the network parameters derived in the training procedure.
@@
Distributions of the total pion energy predicted in training and validation dataset by the three models are compared in Fig. 5 for 100 GeV pions. The energy distributions obtained using the validation dataset do not show any bias when compared to those obtained using the event sample used for training the network. In the bins of the true energy of the charged pions, these distributions are fitted with a Gaussian function in the range of one standard deviation around the mean measured energy. The energy resolution is then obtained as the ratio of the standard deviation ($\sigma$) to the mean ($\mu$) obtained from the fitted Gaussian function. The energy response of pion showers is defined as the $\mu$, normalized to the true energy. The response and resolution as a function of pion incident energy found with the validation dataset matches well with those obtained in the training dataset, as shown in Fig. 6 for the DRN(E) (left), DRN(E,z) (middle), and DRN(E,x,y,z) (right). For the remaining studies, we present the results obtained using the hadron showers in the validation event sample unless stated otherwise.
@@
Figure 6: Comparison of resolution (upper row) and response (bottom row) obtained in training and validation datasets for the models DRN (E) (left), DRN (E,z) (middle), and DRN (E,x,y,z) (right) as function of incident pion energy.
Distributions of energy reconstructed by the three models are shown in Fig. 7 for pions with different energies. We observe that the results obtained restricting to energy information alone are improved when the longitudinal and transverse positions of the rechits are included as input features to the model. The resolution as a function of the pion energy obtained from the three models and for comparison the WS method, are shown in Fig. 8 (top, left). Figure 8 (top, right) shows ratio of resolution obtained using DRN(E), DRN(E,z), and WS method to DRN(E,x,y,z). There is more than a factor of two improvement in the resolution of energy reconstructed using DRN (E) and an additional 20% improvement when the longitudinal information of rechits is added. Adding transverse position information further improves the resolution by up to 10%. These sequential improvements can be interpreted as an indication that the network has learnt how the shower spreads in the calorimeter, and is able to partially correct for the energy leakage of the detector. The energy response obtained using these models is close to true energy of pions as shown in Fig. 8 (bottom).
@@
The large improvement in resolution with DRN (E) over the WS method is attributed to the fact that the information of the spatial pattern of energy deposited is encoded in the distribution of rechit energies obtained using detector level calibration. Distributions of rechit energies measured in units of MIPs in CE-E, CE-H and AHCAL sections, and their distribution in depth in the detector ($z$ coordinates) are shown in Fig. 9. Comparing the corresponding distributions shown in Fig. 4, the rechit energies in units of GeV have clusters with much lower energy in CE-E than in the other two sections. This is also visible from the much larger MIP-to-GeV conversion for electromagnetic section.
@@
Figure 10: Resolution curves obtained with the models DRN(E), DRN(E[MIPs]), DRN(E[MIPs],Flag). By including a flag indicating in which subsection of the calorimeter the rechit belongs to, the resolution obtained with DRN(E) is recovered.

### v3

@@ 2 The HGCAL Prototype, event reconstruction and simulation
The HGCAL prototype detector used in the beam test in 2018 consisted of a silicon sampling electromagnetic section (CE-E prototype) and a silicon sampling hadronic section (CE-H prototype). These were followed by the CALICE AHCAL prototype, which is a sampling calorimeter using scintillator tiles directly read out by SiPMs [7]. The experimental setup is described in [9] and is shown in Fig. 1. The basic detector unit of the CE-E and CE-H prototypes was a silicon module consisting of a 6-inch hexagonal silicon sensor and a printed circuit board (PCB) with the readout electronics, along with a baseplate for mechanical support. Each silicon sensor was subdivided into 128 hexagonal cells with an area of approximately 1.1 $cm^{2}$, and each cell was readout independently with signal amplitude and timing information being collected.
@@
For each event, the signal amplitude and timing information for each detector cell are saved as raw data for offline reconstruction of energies deposited and time of arrival of the signal [11]. The digitized signal from each readout channel, after subtracting pedestal noise, was converted into the corresponding energies measured in the unit of number of minimum ionizing particles (MIPs). The MIP conversion factor for each channel is obtained by fitting the digitized signal obtained with 200 GeV muons with a Landau distribution convoluted with a Gaussian distribution. The maxima of the fitted function is the MIP calibration value for that cell.
The detector setup, along with the beamline elements, was simulated using a Geant4 simulation [12]. The beamline elements, starting from the production target T2 up to the front of the HGCAL prototype, were simulated using the G4Beamline simulation framework implemented in Geant4 version 10.3. The calorimeter sections were simulated using Geant4 version 10.4.3. A schematic of the simulated HGCAL and AHCAL prototype detectors used in the beam test experiment is shown in Fig. 2. The details are summarized in Ref. [11]. The simulated signal is also converted into units of the number of MIPs using the same procedure as used for the data. The position of cells is defined in a right-handed Cartesian coordinate system with the z-axis parallel to the beam direction.
Figure 2: Simulated geometry of the CE-E, CE-H, and AHCAL detector prototypes depicting the position of active layers and absorbing material.
A sample of five million charged pions with a flat energy distribution between 10 and 350 GeV was produced using the Geant4 toolkit implemented in the CMS software framework with FTFP_BERT_EMN physics list [15]. This sample is used to train the neural network models investigated for the energy reconstruction of hadron showers produced by charged pions.
@@ 3 Machine learning strategy and the dynamic reduction network
The highly granular nature of the HGCAL makes it very suitable for the application of machine learning (ML) techniques to reconstruct particle showers in the calorimeter. Compared to other calorimeter designs, the level of detailed information in HGCAL is significantly higher. Machine learning techniques, which can process all or a substantial portion of this data, are particularly effective at leveraging these details to accurately reconstruct hadronic and electromagnetic showers.
@@
These showers have large event-by-event variations in the spatial distributions of energy deposits in the detector cells and are usually interspersed with channels without a measurable signal. Traditional ML architectures like multilayer perceptron neural networks (MLPs) [16] and boosted decision trees (BDTs), expect fixed-size human-designed input features for reconstructing these showers, and it is difficult to design such features to efficiently capture these variations. Modern ML algorithms, such as convolutional neural networks (CNNs), possess powerful feature-learning capabilities, and have demonstrated their ability to outperform traditional ML algorithms, which use human-designed input features, particularly when provided with low level detector information, like calorimeter hits [17, 18, 19, 20, 21].
While image based approaches like CNNs do use low level detector information, the requirement of fixed size inputs necessitate zero-padding or truncation of the input collection. Additionally it is not possible to naturally represent information from multiple layers of detectors with varying granularity in square or rectangular images. Finally, these architectures tend to impose ordering on their inputs that is inappropriate for this permutation-invariant collection of calorimeter hits. A natural way to represent these sparse, high-dimensional, variable-geometry sets of input features is with graphs. In particular, we represent each pion shower as a point cloud of hits in a high-dimensional latent space and then create graphs by drawing edges between nearby hits. Information is allowed to flow along the graph edges in GNNs as in CNNs, but with the advantage of, among others, flexibility in morphology.
@@
In electromagnetic showers, bremsstrahlung and pair-production processes dominate the energy loss until the shower components’ energy falls below the critical energy, typically of the order of 10 MeV. The fluctuations in the amount of energy that is undetected are small and do not affect significantly the energy resolution. However, in hadronic showers a large component of the energy is lost through excitation or disintegration of the absorber nuclei, and, since it is generally undetected, it is known as the ‘invisible energy’ and can amount to up to 40% of the total energy of a shower. Additionally, charged and neutral pions produced in nuclear interactions leave different energy depositions as neutral pions almost instantaneously decay to a pair of photons giving rise to electromagnetic showers while charged pions develop into hadronic showers. Fluctuations in both the production rate of neutral or charged pions, and in the energy lost as ‘invisible energy’ lead to significant event-by-event variations in the energy response of a calorimeter to incoming hadrons.
@@
In the conventional WS method used to determine the energy of an incoming hadron in a calorimeter that is divided into a few sampling layers (for example the 4 or 5 layers of the current CMS hadron calorimeter [22]), where the first layer is usually an electromagnetic calorimeter, the energy deposition in each layer is added with a different weight that is determined either empirically or by simulation. This method does not fully account for the large event-by-event fluctuations in hadronic showers and requires fine tuning for different regions of the detector, as well as for any dependence on the energy of incident hadrons. This method works well for electromagnetic showers, which tend to be uniform, whereas for hadronic showers fluctuations limit its efficacy [23]. In the context of high granularity calorimeters, the method has been extended by using local energy densities as a measure of the electromagnetic component of hadronic showers as described in [24].
@@
The HGCAL, with its fine longitudinal and transverse granularity, produces a highly detailed image of the longitudinal and lateral development of the shower. By using ML methods the information in this three-dimensional image can be used in a structured way to derive the energy and other properties of the incoming hadron. In addition, in the HGCAL, the arrival time of the signal in each readout channel will be available, adding another dimension to the shower reconstruction. However, the potential improvements expected by including timing information for event reconstruction have not been explored in this work.
@@
The DRN [14] builds upon dynamic graph convolution networks [25] with two modifications: firstly the graph operations are performed in a higher dimensional latent space, and secondly, there is an additional graph clustering and pooling step used to iteratively accumulate information across the graph. A schematic of such a network used for the studies presented in this paper is shown in Fig. 3. A summary of the information flow from the inputs – the spatial coordinates of the hits and their energy measured in each detector readout channel – to the required target is as follows.
@@
    The input features, attributes of detector channels such as spatial coordinates (x, y, and z) or energy (E), are normalized using fixed values such that most of the values lie in the range from zero to one, with outliers allowed. The normalized input data are mapped onto a 64-dimensional latent space using a fully connected neural network, which is an MLP encoder with a depth of three layers that are connected via an exponential linear unit activation function.
@@
    High-level features are learned using GNN techniques by repeating the following three steps twice:
@@
      Graph convolution step: the resulting graphs are processed using the "EdgeConv" operation [26], where two-layer fully connected networks are used to create "messages" that are sent along the edges of the graph. These "messages" are aggregated at each node in the graph to form the new node features.
@@
      Graph clustering and pooling: the graph convolution step is followed by graph clustering using the graclus clustering algorithm [27] and a pooling step, in which the updated features are weighted by distance and the clusters are combined using a max-pooling strategy to form a single point. The graclus clustering algorithm successively and pairwise, combines connected graph nodes within a cluster that maximizes the resulting edge weight of the combined node.
    Each pass through these three steps – graph generation, graph convolutions, and graph clustering and pooling – halves the number of nodes in the graph.
@@
As described in Ref. [14], the clustering and pooling steps are optimized in a way to maximize the reduction in information passed on for further processing, hence the network is called a "dynamic reduction network (DRN)".
Figure 3: This schematic illustrates the GNN-based dynamic reduction network. The $x_{i}$ represents rechits, with red edges showing their connections in the latent space which are updated during graph convolution (in magenta). In the graph clustering stage, small circles indicate clusters, and nodes within as blue dots. Final predictions are produced after passing through the output neural network.
@@ 4 Energy Reconstruction of pion showers using DRN method
A sample of simulated hadron showers produced by negatively charged pions, uniformly sampled in the energy range from 10 to 350 GeV, is used to train the DRN for the reconstruction of the energy of incident pions. The hadron showers were simulated using the same Geant4 geometry and physics lists that were used for the simulation of events in the beam tests described above. We apply the same selection criteria to remove events where the particles loose a large fraction of their energy in the beamline and apply the same noise rejection thresholds that were used in the data [11]. The input to the DRN is a point cloud of the detector channels associated with a given hadronic shower. The attributes of each channel, used for training the model, are its position coordinates, $x$, $y$, and $z$, given in centimeters and the amplitude of the energy deposition in the sensor as either the number of MIPs or in terms of the effective energy in GeV, defined below.
As this is a sampling calorimeter, we need to infer from the signal in each sensor the energy deposited in the absorbers. In the prototype detector, the absorbers in CE-E, CE-H, and AHCAL were different and hence energy scale factors need to be estimated separately for them. We estimate the energy deposited in the absorber that corresponds to the observed signal amplitude using a MIP-to-GeV conversion factor. For CE-E this factor is estimated with 50 GeV positrons. In CE-H and AHCAL the MIP-to-GeV conversion factor is estimated with 50 GeV pions that do not initiate a hadronic shower in CE-E. We obtain factors of approximately 10.5 MeV per MIP and 80 MeV per MIP for the electromagnetic and hadronic sections, respectively.
The signals in the detector cells, weighted by their corresponding MIP-to-GeV factor, are used to estimate the amount of energy deposited in the absorber plates. These estimates of energy deposited in the absorber together with their positions are referred to as "reconstructed hits" or "rechits". Distributions of the number of rechits with a signal measured above the noise thresholds in CE-E, CE-H, and AHCAL sections, and their $x$ and $y$ coordinates are shown in Fig. 4 (top left) and Fig. 4 (top right), respectively, for simulated 200 GeV $\pi^{-}$ data. The distributions of the energy estimated in units of GeV in the detector cells of the three detector sections are shown in Fig. 4 (bottom left). Fig. 4 (bottom right) shows the distribution of rechit energies along the depth of the detector (z coordinate).
@@
The energies of all rechits can be summed to obtain an estimate of the total energy of the particles; this sum is referred to as $E_{fix}$ in the following. However, since the response of a sampling calorimeter, such as the HGCAL prototype, to hadrons is not linear with the incident particle’s energy, $E_{fix}$ is a suboptimal estimate of the particle’s true energy. The WS method uses a $\chi^{2}-minimization$ method for a given incident particle energy, as discussed in [11], to obtain weights to combine energies measured in CE-E, CE-H, and AHCAL sections. Although the linearity of the response can be restored and the resolution improved, this method does not account for event-by-event fluctuations in the hadronic showers.
@@
We have trained a DRN model with simulated hadronic showers generated with a flat incident energy distribution. The target in the ML training was the ratio of the true energy of the particle to that measured using the fixed MIP-to-GeV weights ($E_{fix}$). In the training, the learning rate was set to 0.0001. We trained three different models with three different input features, the energy ($E$), the energy and the longitudinal position ($E,z$), and the energy and both the lateral and the longitudinal positions ($E,x,y,z$) of each detector channel. These models are trained until the value of the loss function evaluated on the validation dataset stops improving in subsequent epochs. The epoch numbers 110, 114, and 135 correspond to the minimum loss values for DRN(E), DRN(E,z), and DRN(E,x,y,z), respectively, and corresponding training weights are used for evaluating the performance of pion energy reconstruction in the following. Approximately 20% of the generated events were reserved in a separate dataset for validation of the models using the network parameters derived in the training procedure. It took approximately 25 minutes to train a model for a single epoch on 4 NVIDIA V100 GPUs, with 192 GB RAM on each of the 4 GPU nodes, in parallel [28].
@@
Distributions of the total pion energy predicted in training and validation dataset by the three models are compared in Fig. 5 for 100 GeV pions. The energy distributions obtained using the validation dataset do not show any bias when compared to those obtained using the event sample used for training the network. In the bins of the true energy of the charged pions, these distributions are fitted with a Gaussian function in the range of one standard deviation around the mean measured energy. The energy resolution is then obtained as the ratio of the standard deviation ($\sigma$) to the mean ($\mu$) obtained from the fitted Gaussian function. The energy response of pion showers is defined as the $\mu$, normalized to the true energy. The response and resolution as a function of pion incident energy found with the validation dataset match well with those obtained in the training dataset, as shown in Fig. 6 for the DRN(E) (left), DRN(E,z) (middle), and DRN(E,x,y,z) (right). For the remaining studies, we present the results obtained using the hadron showers in the validation event sample unless stated otherwise.
@@
Figure 6: Comparison of resolution (upper row) and response (bottom row) obtained in training and validation datasets for the models DRN (E) (left), DRN (E,z) (middle), and DRN (E,x,y,z) (right) as a function of incident pion energy.
Distributions of energy reconstructed by the three models are shown in Fig. 7 for pions with different energies. We observe that the results obtained restricting to energy information alone are improved when the longitudinal and transverse positions of the rechits are included as input features to the model. The resolution as a function of the pion energy obtained from the three models and for comparison the WS method are shown in Fig. 8 (top, left). Figure 8 (top, right) shows the ratio of resolution obtained using DRN(E), DRN(E,z), and WS method to DRN(E,x,y,z). There is more than a factor of two improvement in the resolution of energy reconstructed using DRN (E) and an additional 20% improvement when the longitudinal information of rechits is added. Adding transverse position information further improves the resolution by up to 10%. These sequential improvements can be interpreted as an indication that the network has learned how the shower spreads in the calorimeter, and is able to partially correct for the energy leakage of the detector. The energy response obtained using these models is close to the true energy of pions as shown in Fig. 8 (bottom).
@@
The large improvement in resolution with DRN (E) over the WS method is attributed to the fact that the information of the spatial pattern of energy deposited is encoded in the distribution of rechit energies obtained using detector level calibration. Distributions of rechit energies measured in units of MIPs in CE-E, CE-H, and AHCAL sections, and their distribution in depth in the detector ($z$ coordinates) are shown in Fig. 9. Comparing the corresponding distributions shown in Fig. 4, the rechit energies in units of GeV have clusters with much lower energy in CE-E than in the other two sections. This is also visible from the much larger MIP-to-GeV conversion for electromagnetic section.
@@
Figure 10: Resolution curves obtained with the models DRN(E), DRN(E[MIPs]), DRN(E[MIPs],Flag). By including a flag indicating in which subsection of the calorimeter the rechit belongs, the resolution obtained with DRN(E) is recovered.